OpenStudy (anonymous):

find all real numbers that satisfy the equation |1-3x|=3

OpenStudy (anonymous):

sin is negative in the third and fourth quadrants in the third quadrant, sin 240--sqrt(3)/2 in the fourth quadrant, sin 300=-sqrt(3)/2, so the angles are 240 and 300 deg, or 4pi/3 and 5 pi/3 radians you can have an infinite number of answers to this, if you remember that sin is 2 pi periodic, so your complete answer set is 4 pi /3 + 2 n pi 5 pi/3 + 2 n pi where n is an integer

OpenStudy (anonymous):

@abayomi12 what are u doing?

OpenStudy (anonymous):

the answer is -2/3 and 4/3, but i dont know how to get there, your methode cis not quite those values ar least not that i can interpert calculus

OpenStudy (anonymous):

I1-3xI=3 SO, 1-3x=3 and 1-3x=-3

OpenStudy (anonymous):

Can u solve that two equations

OpenStudy (anonymous):

sauravshakya have a look back at the answer i gave a moment ago

OpenStudy (anonymous):

no

OpenStudy (anonymous):

1-3x=3 and 1-3x=-3

OpenStudy (anonymous):

WE need to solve that

OpenStudy (anonymous):

am doing thing

OpenStudy (anonymous):

1-3x=3 3x=1-3 3x=-2 x=-2/3

OpenStudy (anonymous):

which is one solution

OpenStudy (anonymous):

my answer is the same with yours just the method is different

OpenStudy (anonymous):

now, solve 1-3x=-3

OpenStudy (anonymous):

to get another answer

OpenStudy (anonymous):

sauravshakya im a fan

OpenStudy (anonymous):

@lgbasallote did u understand what @abayomi12 did

OpenStudy (anonymous):

no

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

can u get another solution @sparksaflyen

OpenStudy (anonymous):

no

OpenStudy (anonymous):

i will for that

OpenStudy (anonymous):

solve 1-3x=-3

OpenStudy (anonymous):

can u?

OpenStudy (anonymous):

a minute trying to get another method

OpenStudy (anonymous):

yeah i will but trying to get another

OpenStudy (lgbasallote):

the results werent the same

OpenStudy (anonymous):

thats right thanks, obviously its because im working with absolute values

OpenStudy (lgbasallote):

btw i was referring to wha @sauravshakya said to me

OpenStudy (anonymous):

First you need to distribute (multiply 1/3 by x and -3). This will give you 1/3x-1<(x-2)/4 I'm assuming that x-2 is all over 4 on the right side. Then you'll have to multply